OpenStudy (jollypikachu300):

15/8 divided by 9/10 ? a-25/36 b-9/16 c-45/80 d-1/36

pooja195 (pooja195):

\[\huge~\rm~\frac{ 15\times 10=? }{ 8\times 9=?}\]

OpenStudy (photon336):

yeah that's how I would do it. Keep change the sign from division to multiplication and flip the second number. then multiply across the numerator and denominator.

OpenStudy (jollypikachu300):

150/72

pooja195 (pooja195):

Good now simplify that

OpenStudy (jollypikachu300):

how do i simplify

pooja195 (pooja195):

can you find the GCF of 150 and 72?

OpenStudy (jollypikachu300):

no

pooja195 (pooja195):

We know they are both even numbers right so 2 is possible 150 divided by 2=? 72 divided by 2 =?

OpenStudy (jollypikachu300):

75/36

OpenStudy (jollypikachu300):

no 25/36

OpenStudy (jollypikachu300):

right

pooja195 (pooja195):

The answer im getting isnt one of your answer choices...did you type the problem wrong?

OpenStudy (jollypikachu300):

idk

pooja195 (pooja195):

"idk" How about you check the problem again to double check that u didnt make a mistake?

OpenStudy (jollypikachu300):

cant u tell me the anwser

OpenStudy (whpalmer4):

It may be easier to see how to simplify this if you factor the numbers before multiplying them, then cancel terms which appear in both numerator and denominator: \[\frac{15}{8}\div\frac{9}{10} = \frac{15}{8}*\frac{10}{9} = \frac{15*10}{8*9} = \frac{3*5*2*5}{2*2*2*3*3} \]\[= \frac{\cancel{2*}\cancel{3*}5*5}{\cancel{2*}\cancel{3*}2*2*3}=\frac{5*5}{2*2*3} = \] I agree that the correct answer to the question as you have provided it here does not appear in your list of answer choices.